Six identical capacitors with capacitance C are co


1)What is the potential difference between points a and b?

2) What is the equivalent capacitance of these six capacitors?
Answer #1
Concepts and reason

The concepts used to solve this problem are series and parallel combination of capacitance, and the charge across the capacitance.

Firstly, calculate the equivalent capacitance of the circuit by using the series and parallel combination of the capacitance, after that calculate the charge across the circuit by using the expression of charge in term of capacitance and voltage.

Fundamentals

Capacitance is the ability of a conductor to store an electric charge. The SI unit of capacitance is Farad.

If NN number of capacitors are connected in parallel then the equivalent capacitance is, Ceq=C1+C2+.......+CN{C_{{\rm{eq}}}} = {C_1} + {C_2} + ....... + {C_{\rm{N}}}

If NN number of capacitors are in series then the equivalent capacitance is given by:

1Ceq=1C1+1C2.....+1CN\frac{1}{{{C_{eq}}}} = \frac{1}{{{C_1}}} + \frac{1}{{{C_2}}}..... + \frac{1}{{{C_N}}}

The voltage on the capacitor is given by the voltage division rule:

VX=VSCTCX{V_X} = {V_S}\frac{{{C_T}}}{{{C_X}}}

Here, CT{C_T} is the total capacitance and VS{V_S} is the source voltage.

The equivalent circuit for the circuit given is as follows,

C
C)
12 V
C
b
a
C
C
Part A

12 V
C.
C
C.
2
2
2
Part B

3C
12 V
2
Part C

(a)

The capacitors are same at point aa and bb .The voltage at aa is the voltage dropped due to one capacitor CC , as the voltage source has one capacitor CC between it and point aa .

Va=V{V_a} = V ……(1)

Similarly, the voltage is dropped, when it reaches bb , by the same amount as aa . Again there is one capacitor CC between the source and point bb .

Vb=V{V_b} = V ……(2)

The potential at point aa is the same as points bb so the potential difference between points a and b is,

VaVb=0ΔV=0\begin{array}{c}\\{V_a} - {V_b} = 0\\\\\Delta V = 0\\\end{array}

[Part a]

Part a

Part a

(b)

If NN number of capacitors are in series then the equivalent capacitance is given by:

1Ceq=1C1+1C2.....+1CN\frac{1}{{{C_{eq}}}} = \frac{1}{{{C_1}}} + \frac{1}{{{C_2}}}..... + \frac{1}{{{C_N}}}

From the diagram part A, two capacitors are connected in series:

1Cequ=1C+1C=2CCequ=C2\begin{array}{c}\\\frac{1}{{{C_{equ}}}} = \frac{1}{C} + \frac{1}{C}\\\\ = \frac{2}{C}\\\\{C_{equ}} = \frac{C}{2}\\\end{array}

If NN number of capacitors are connected in parallel then the equivalent capacitance is, Ceq=C1+C2+.......+CN{C_{{\rm{eq}}}} = {C_1} + {C_2} + ....... + {C_{\rm{N}}}

Form the diagram part B, three capacitors are connected in parallel.

Cequ=C2+C2+C2=3C2\begin{array}{c}\\{C_{equ}} = \frac{C}{2} + \frac{C}{2} + \frac{C}{2}\\\\ = \frac{{3C}}{2}\\\end{array}

[Part b]

Part b

Part b

Ans: Part a

The potential difference between points aa and bb is 0V0{\rm{ V}} .

Part b

The equivalent capacitance of the six capacitors is 3C2\frac{{3C}}{2} .